OpenStudy (stamp):

[CALCULUS III—CHAIN RULE] Find ∂w/∂s and ∂w/∂t where w = y^3 −3 x^2 y, x = e^s, y = e^t

OpenStudy (stamp):

\[w=y^3-3x^2y\]\[x=e^s,\ y=e^t\]

OpenStudy (phi):

\[ \frac{d}{ds} w = \frac{d}{ds} y^3 - 3 \frac{d}{ds} (x^2y) \]

OpenStudy (phi):

\[= 3 y^2 \frac{dy}{ds} - 3 x^2\frac{dy}{ds}- 6 x y \frac{dx}{ds} \] now find dy/ds and dx/ds

OpenStudy (stamp):

\[y=e^t,\ dy/ds = 0\]\[x=e^s,\ dx/ds=e^s\]

OpenStudy (phi):

looks good. \[ \frac{dw}{ds} = -6xye^s \]

OpenStudy (phi):

sub in for x and y to get things in terms of s and t

OpenStudy (stamp):

^ This is the partial ? If so we repeat a similar process for ∂w/∂t

OpenStudy (phi):

yes, repeat for dw/dt

OpenStudy (stamp):

\[∂w/∂s=-6xye^s=-6e^{2s}e^t\]

OpenStudy (phi):

or 2s+t for the exponent

OpenStudy (stamp):

^ sounds better, I did not think to combine the exponents. Very nice.